OpenStudy (anonymous):

Write the equation of the line that passes through the point 4 comma 1 and has a slope of 2 in point-slope form.

OpenStudy (anonymous):

point slope form is (y-y0)=m(x-x0) we have (x0,y0) as (4,1) and m=2 thus we have (y-1)=2(x-4) y-1=2x-8 2x-8+1-y=0 2x-y-7=0